OpenStudy (anonymous):

passing through (-1,3) and (4,3)

OpenStudy (accessdenied):

What is passing through them? A line? A curve? A car? Ghosts...?

OpenStudy (anonymous):

If you can give us/me some more details, I can probably can help you with this one, I love graphs.

OpenStudy (anonymous):

They are asking us to find the function that passes through those two points. @FeoNeo33 This is actually enough information.

OpenStudy (blockcolder):

Your best bet is a quadratic function. Let our mystery function f(x):=ax^2+bx+c. We solve the system \[\begin{cases} 3=a+b+c \\ 3=16a+4b+c \end{cases}\] We find that a, b, and c need to satisfy \[15a+3b=0\\ \Rightarrow 5a+b=0\\ \Rightarrow b=-5a\] Now if we had one more point we could figure out precisely what a, b, and c are.

OpenStudy (accessdenied):

There's not enough information to determine what type of function exactly. Like, you could have a basic linear function y=3, or maybe its a set of quadratic functions like determined above. He'd have to specify what we were looking for to make an accurate answer. It's sometimes dangerous to assume things. :P